OpenStudy (anonymous):

Algebra 2/Trigonometry help, please? I have to prove cos θ - cos θ•sin^2 θ = cos^3 θ.

OpenStudy (anonymous):

\[\cos \theta -\cos \theta \times \sin ^{2}\theta =\cos ^{3}\theta \]

OpenStudy (anonymous):

\[Cosx (1-\sin ^{2}x )= cosx *\cos ^{2}x\]

OpenStudy (anonymous):

Thank you very much.

OpenStudy (anonymous):

You are welcome :)